# APCoTTA: Continual Test-Time Adaptation for Semantic Segmentation of Airborne LiDAR Point Clouds

### v1

@@ 4.2 Baselines
We compare our method with multiple CTTA baselines, including Source, BN Stats Adapt [40], Pseudo-label [23], TENT-online [51], TENT-Continual [51], CoTTA [53], and PALM [28].
For Source, we apply the pretrained model in evaluation mode on the benchmark. BN Stats Adapt predicts using batch normalization statistics from the current batch without modifying network weights. Pseudo-labeling generates pseudo-labels based on the model’s highest-confidence predictions and incorporates them as ground truth for training. TENT minimizes the model’s prediction entropy during testing, updating feature normalization statistics and optimizing channel-wise affine transformation parameters. TENT-online resets the model to its pretrained state when encountering a new domain. TENT-Continual continuously adapts to new domains without resetting the model. CoTTA employs weight averaging and augmentation-averaged pseudo labels while randomly restoring a portion of the weights to the source-pretrained state for online adaptation. PALM automatically selects layers for adaptation based on model prediction uncertainty and adjusts the learning rate according to parameter sensitivity, enabling continual test-time adaptation.
@@ 4.3 Implementation Details
We evaluate our approach on two continuous test adaptation tasks: ISPRS-to-ISPRSC and H3D-to-H3DC.
For ISPRS-to-ISPRSC and H3D-to-H3DC, We train the network on original ISPRS or H3D training data and then sequentially adapt the pretrained network to each corruption type. Following [53], we evaluate all models under the largest corruption severity level 5. Following [52], We use KPConv [47] as the backbone. Due to the limitations of GPU memory, for the ISPRSC and H3DC datasets, the overlapping spherical sub-cloud radii are 20 m and 11 m, and the batch sizes are 2 and 8. All other parameters are kept consistent with the default settings of the KPConv segmentation network. We adopt the same Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.98 and an initial learning rate of $10^{-2}$. For both datasets, the threshold $S_{0}$ for the DSTL module, $\tau$ for the EGCL module, and $\alpha$ for the RPI module are set to 0.001, 0.8, and 0.999, respectively. All models are implemented in PyTorch 1.8.2 and trained on a single GeForce GTX 3090 24GB GPU.
@@
[…]
Method scene space uniform density cutout impulse gaussian Mean mIoU Mean OA
@@
[…]
Method scene space uniform density cutout impulse gaussian Mean mIoU Mean OA
@@ 4.5.1 ISPRS to ISPRSC
To validate the effectiveness of our method, we compare it with several existing approaches on the continual test-time semantic segmentation task from ISPRS to ISPRSC. As shown in Tab. 1, directly using the non-adapted pretrained model results in a low mIoU of 40.82%, mainly due to the significant domain shift between source and target domains. The global mean and variance from the source domain fail to generalize to the target domain, and the performance is further degraded by catastrophic forgetting and error accumulation. The BN Stats Adapt method, which updates only the batch-wise mean and variance while keeping other parameters fixed, improves performance by 4.33% in mIoU. In contrast, pseudo-labeling methods lead to mIoU drops, as incorrect pseudo-labels introduce cumulative errors and misleading signals, reducing the model’s adaptability. TENT-online leverages prior knowledge of domain shifts, resetting the model to its pretrained state upon domain change, which further enhances adaptation. However, this assumption is impractical, as domain change information is typically unavailable in real scenarios. TENT-continual adapts continuously without resetting the model and achieves an improvement of about 5% in mIoU, though it still underperforms compared to the online version, indicating persistent issues with forgetting and error accumulation. CoTTA adopts a Mean Teacher model to refine pseudo-labels and periodically resets model weights, achieving significant gains. PALM preserves most of the pretrained knowledge and adjusts the learning rate dynamically, resulting in a 5.34% improvement in mIoU. Compared to all these methods, our approach achieves the best performance across multiple target domains, with an mIoU improvement of approximately 9%, as highlighted in bold in Tab. 1. By losing out low-confidence samples, we effectively mitigate error accumulation. Additionally, we randomly select a subset of parameters to be weighted with the pretrained model, balancing target adaptation and the retention of source knowledge, thereby reducing catastrophic forgetting.
@@ 4.5.2 H3D to H3DC
To further demonstrate the effectiveness of our approach, we evaluated it on a more challenging continual test-time adaptation task, H3D to H3DC. As shown in Tab. 2, our method achieves the best performance among all compared state-of-the-art approaches, with a significant mIoU improvement of 13.8%. While methods such as BN Stats Adapt, Pseudo-label, TENT-online, TENT-continual, CoTTA, and PALM achieved around 8% mIoU improvement, they suffer from severe catastrophic forgetting and error accumulation over prolonged adaptation, leading to poor performance on target domains with impulse and Gaussian noise. In contrast, our method effectively mitigates catastrophic forgetting and error accumulation, resulting in significantly better adaptation in noisy target domains. Additionally, as illustrated in Fig. 7, our method demonstrates superior class-wise adaptation across multiple categories, whereas other methods exhibit notable class misclassification. More detailed results are in the .2.
@@ 4.6.1 Effectiveness of Each Component
We conduct ablation studies on two continual test-time adaptation tasks, ISPRS to ISPRSC and H3D to H3DC, to evaluate the effectiveness of each module. As shown in Tab. 3, when all modules are removed, the model adapts to the target domain using only the current batch’s mean and variance, a fully learnable shared-weight network, and a consistency loss. Due to catastrophic forgetting and error accumulation, the model performs poorly, with mIoU of 25.83% and 33.37% on the two tasks, respectively. When the first module, DSTL, is enabled, the model retains most of the useful knowledge from the source domain, which effectively mitigates catastrophic forgetting and significantly improves performance. Enabling the second module, EBCL, further reduces the influence of erroneous information by losing inaccurate samples, thereby alleviating error accumulation. Finally, when the third module, RPI, is activated, the model maintains strong adaptation to the target domain while preserving alignment with the source domain distribution, further mitigating catastrophic forgetting.
@@ 4.6.2 Hyperparameter Sensitivity
We analyze the sensitivity of our model to three key hyperparameters in two continuous test-time adaptation tasks (ISPRS to ISPRSC and H3D to H3DC), as shown in Fig. 8: Select layer threshold $S_{0}$: This parameter controls the number of trainable layers. We vary it within the range 0, 0.0005, 0.001, 0.0015, 0.002. The best performance is achieved around 0.001, with noticeable performance drops when deviating from this value. Entropy threshold $\tau$: This parameter filters out samples with very low confidence. We test values in the range 0.7, 0.75, 0.8, 0.85, 0.9. The results show that model performance is relatively stable across this range. Weighting coefficient $\alpha$: This parameter regulates the distance between the adaptive model and the source domain. Our method achieves optimal performance when $\alpha$ is set to 0.999.
@@ 5 Conclusions
In this study, we first address the lack of continuous test-time adaptation (CTTA) benchmarks for airborne LiDAR point cloud semantic segmentation by introducing two new datasets: ISPRSC and H3DC. To tackle the performance degradation caused by domain shifts in real-world scenarios, we propose the first CTTA method specifically designed for airborne LiDAR point clouds, named APCoTTA. It comprises three key components: (1) a dynamic trainable layer selection strategy, which freezes most network layers while updating only a few crucial ones, mitigating catastrophic forgetting and preserving source domain knowledge; (2) an entropy-based consistency loss that filters out low-confidence samples to reduce error accumulation; and (3) a random parameter interpolation mechanism that blends the parameters of trainable layers with their source counterparts to further alleviate forgetting. Evaluated on our proposed benchmarks, APCoTTA achieves mIoU improvements of approximately 9% and 14% over direct inference.
Despite the progress made in this study, several limitations remain. The issue of class imbalance in airborne LiDAR point clouds, which can significantly impact model performance on long-tail or rare classes, has not been thoroughly addressed. Future work may explore more sophisticated re-sampling strategies or class-balanced weighting methods to tackle this challenge. Furthermore, we plan to extend continuous test-time adaptation to open-set scenarios to handle the emergence of unknown classes during deployment.
@@ Acknowledgements
This work was supported by the National Key Research and Development Program of China (2024YFF0810400), and the National Natural Science Foundation of China under Grant 42201481, U22A20556).

### v2

@@ 4.2 Baselines
We compare our method with multiple CTTA baselines, including Source, BN Stats Adapt [20], Pseudo-label [18], TENT-online [41], TENT-continual [41], CoTTA [43], PALM [23], and Wang et al. [42].
For Source, we apply the pretrained model in evaluation mode on the benchmark. BN Stats Adapt [20] predicts using batch normalization statistics from the current batch without modifying network weights. Pseudo-label [18] generates pseudo-labels based on the model’s highest-confidence predictions and incorporates them as ground truth for training. TENT [41] minimizes the model’s prediction entropy during testing, updating feature normalization statistics and optimizing channel-wise affine transformation parameters. TENT-online resets the model to its pretrained state when encountering a new domain. TENT-continual continuously adapts to new domains without resetting the model. CoTTA [43] employs weight averaging and augmentation-averaged pseudo labels while randomly restoring a portion of the weights to the source-pretrained state for online adaptation. PALM [23] automatically selects layers for adaptation based on model prediction uncertainty and adjusts the learning rate according to parameter sensitivity, enabling continual test-time adaptation. Wang et al. [42] employ a Progressive Batch Normalization (PBN) module that updates BN statistics using the exponential moving average of test batches, while concurrently optimizing the affine parameters of BN layers via a self-supervised loss function. Specifically dsigned for TTA, this method assumes a static target domain, whereas CTTA addresses ongoing domain shifts.
@@ 4.3 Implementation Details
We evaluate our approach on two continuous test adaptation tasks: ISPRS-to-ISPRSC and H3D-to-H3DC. Specifically, we first pre-train the source models on the training sets of the original ISPRS and H3D datasets. Then, we sequentially adapt the pre-trained network to the corruption benchmarks (ISPRSC and H3DC), which are constructed from the testing set of ISPRS and the validation set of H3D, respectively. Note that during adaptation, the ground truth labels of the target domain are strictly withheld and used only for performance evaluation. Following [43], we evaluate all models under the largest corruption severity level 5. Following [42], We use KPConv [37] as the backbone. To accommodate GPU memory constraints, for the ISPRSC and H3DC datasets, the overlapping spherical sub-cloud radii are 20 m and 11 m, and the batch sizes are 2 and 8. All other parameters are kept consistent with the default settings of the KPConv segmentation network. We adopt the same Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.98 and an initial learning rate of $10^{-2}$. For both datasets, the threshold $S_{0}$ for the DSTL module, $\tau$ for the EBCL module, and $\alpha$ for the RPI module are set to 0.001, 0.8, and 0.999, respectively. All models are implemented in PyTorch 1.8.2 and trained on a single GeForce GTX 3090 24GB GPU.
@@
Time $t\xrightarrow{\makebox[200.0pt][c]{\quad}}$
Method sunlight space uniform density cutout impulse gaussian Mean mIoU Mean OA
@@
Time $t\xrightarrow{\makebox[200.0pt][c]{\quad}}$
Method sunlight space uniform density cutout impulse gaussian Mean mIoU Mean OA
@@ 4.5.1 ISPRS to ISPRSC
To validate the effectiveness of our method, we compare it with several existing approaches on the continual test-time semantic segmentation task from ISPRS to ISPRSC. As shown in Tab. 1, directly using the non-adapted pretrained model results in a low mIoU of 40.82%, mainly due to the significant domain shift between source and target domains. The global mean and variance from the source domain fail to generalize to the target domain, and the performance is further degraded by catastrophic forgetting and error accumulation. The BN Stats Adapt [20] method, which updates only the batch-wise mean and variance while keeping other parameters fixed, improves performance by 4.33% in mIoU. In contrast, pseudo-labeling methods lead to mIoU drops, as incorrect pseudo-labels introduce cumulative errors and misleading signals, reducing the model’s adaptability. TENT-online[41] leverages prior knowledge of domain shifts, resetting the model to its pretrained state upon domain change, which further enhances adaptation. However, this assumption is impractical, as domain change information is typically unavailable in real scenarios. TENT-continual [41] adapts continuously without resetting the model and achieves an improvement of about 5% in mIoU, though it still underperforms compared to the online version, indicating persistent issues with forgetting and error accumulation. CoTTA [43] adopts a Mean Teacher model to refine pseudo-labels and periodically resets model weights, achieving significant gains. PALM [23] preserves most of the pretrained knowledge and adjusts the learning rate dynamically, resulting in a 5.34% improvement in mIoU. Wang et al. [42] propose a TTA method tailored for geospatial point clouds, achieving a superior mIoU of 47.21%, surpassing previous baselines. However, it still remains inferior to our APCoTTA (49.74%), as its standard TTA design lacks specific mechanisms to handle the continuous domain shifts inherent in the CTTA setting. Compared to all these methods, our approach achieves the best performance across multiple target domains, with an mIoU improvement of approximately 9%, as highlighted in bold in Tab. 1. By filtering out low-confidence samples, we effectively mitigate error accumulation. Additionally, we randomly select a subset of parameters to be weighted with the pretrained model, balancing target adaptation and the retention of source knowledge, thereby reducing catastrophic forgetting.
@@ 4.5.2 H3D to H3DC
To further demonstrate the effectiveness of our approach, we evaluated it on a more challenging continual test-time adaptation task, H3D to H3DC. As shown in Tab. 2, our method achieves the best performance among all compared state-of-the-art approaches, with a significant mIoU improvement of 13.8%. While methods such as BN Stats Adapt [20], Pseudo-label [18], TENT-online[41], TENT-continual [41], CoTTA [43], and PALM [23] achieve around 8% mIoU improvement, they suffer from severe catastrophic forgetting and error accumulation over prolonged adaptation, leading to poor performance on target domains with impulse and Gaussian noise. Wang et al. [42], which is designed for remote sensing point clouds, demonstrates strong performance in the early stages and achieves the second-best overall mIoU of 41.32%. However, its performance degrades noticeably in the later stages of the corruption sequence (e.g., 32.00% on Impulse and 31.52% on Gaussian noise), indicating a susceptibility to error accumulation over long-term adaptation. In contrast, our method effectively mitigates catastrophic forgetting and error accumulation, resulting in significantly better adaptation in target domains. Additionally, as illustrated in Fig. 7, our method demonstrates superior class-wise adaptation across multiple categories, whereas other methods exhibit notable class misclassification. More detailed results are provided in the .2.
@@ 4.6.1 Effectiveness of Each Component
We conduct ablation studies on two continual test-time adaptation tasks, ISPRS to ISPRSC and H3D to H3DC, to evaluate the effectiveness of each module. As shown in Tab. 3, when all modules are removed, the model adapts to the target domain using only the current batch’s mean and variance, a fully learnable shared-weight network, and a consistency loss. Due to catastrophic forgetting and error accumulation, the model performs poorly, with mIoU of 25.83% and 33.37% on the two tasks, respectively. When the first module, DSTL, is enabled, the model retains most of the useful knowledge from the source domain, which effectively mitigates catastrophic forgetting and significantly improves performance. Enabling the second module, EBCL, further reduces the influence of erroneous information by discarding inaccurate samples, thereby alleviating error accumulation. Finally, when the third module, RPI, is activated, the model maintains strong adaptation to the target domain while preserving alignment with the source domain distribution, further mitigating catastrophic forgetting.
@@ 4.6.2 Hyperparameter Sensitivity
We analyze the sensitivity of our model to four key hyperparameters in two continuous test-time adaptation tasks (ISPRS to ISPRSC and H3D to H3DC), as shown in Fig. 8: Select layer threshold $S_{0}$: This parameter controls the number of trainable layers. We vary it within the range 0, 0.0005, 0.001, 0.0015, 0.002. The best performance is achieved around 0.001, with noticeable performance drops when deviating from this value. Entropy threshold $\tau$: This parameter filters out samples with very low confidence. We test values in the range 0.7, 0.75, 0.8, 0.85, 0.9. The results show that model performance is relatively stable across this range. Weighting coefficient $\alpha$: This parameter regulates the distance between the adaptive model and the source domain. Our method achieves optimal performance when $\alpha$ is set to 0.999. Sampling probability $p$: We evaluate $p\in\{0.001,0.005,0.01,0.05,0.1\}$, and results demonstrate that performance peaks at $p=0.01$. Notably, increasing $p$ to $0.1$ causes a significant drop in mIoU (e.g., $3.40\%$ on ISPRSC and $3.12\%$ on H3DC), indicating that excessive restoration overly constrains the model to the source domain, hindering adaptation to extreme geometric shifts. Conversely, lowering $p$ to $0.001$ results in insufficient regularization against forgetting. Thus, $p=0.01$ provides the optimal cumulative regularization effect.
@@ 5 Conclusions
In this study, we present the dedicated benchmarks for CTTA in airborne LiDAR point cloud semantic segmentation, filling a critical gap in the field through the release of the ISPRSC and H3DC datasets. To address the performance degradation caused by continuous domain shifts in real-world scenarios, we propose a CTTA framework specifically designed for airborne LiDAR point clouds, named APCoTTA. It comprises three key components: (1) a dynamic trainable layer selection strategy, which freezes most network layers while updating only a few crucial ones, mitigating catastrophic forgetting and preserving source domain knowledge; (2) an entropy-based consistency loss that filters out low-confidence samples to reduce error accumulation; and (3) a random parameter interpolation mechanism that blends the parameters of trainable layers with their source counterparts to further alleviate forgetting. Evaluated on our proposed benchmarks, APCoTTA achieves mIoU improvements of approximately 9% and 14% over direct inference.
Despite the progress made in this study, several limitations remain. The issue of class imbalance in airborne LiDAR point clouds, which can significantly impact model performance on long-tail or rare classes, has not been thoroughly addressed. Future work may explore more sophisticated re-sampling strategies or class-balanced weighting methods to address this challenge. Furthermore, we plan to extend continuous test-time adaptation to open-set scenarios to handle the emergence of unknown classes during deployment.
@@ Acknowledgements
This work was supported by the National Key Research and Development Program of China (2024YFF0810400), and the National Natural Science Foundation of China under Grant 42201481, U22A20556.